OpenStudy (anonymous):

A ball is drawn from a box containing 5 red, 8 black and 7 blue balls. Determine the probability that: a) if a ball is drawn at random, it is black. answer i got: 0.40 b) if two balls are drawn, they are both black. answer i got: 0.1474 c) if three balls are drawn, exactly one is black, one is blue and one is red. d) if three balls are drawn, at least one is black. e) if four balls are drawn, at most one of them is black. Note: All selection is without replacement. can you please check the answers i got is right or wrong and can you help me solve c, d, and e too.

OpenStudy (anonymous):

a and b are correct

OpenStudy (anonymous):

great :) could you help me solve the other ones as well. i'm keep getting wrong answers i think i'm doing something wrong. no idea

OpenStudy (anonymous):

c) total balls= 20 8/20 * 7/19 * 5/18 = 0.0409

OpenStudy (anonymous):

i got had got the same answer for c

OpenStudy (anonymous):

so is it wrong??

OpenStudy (anonymous):

i wasn't sure if i did it right or wrong i wanted to make sure i did it right, since you got the same answer it should be correct

OpenStudy (anonymous):

d) 8/20 * 7/19 * 6/18

OpenStudy (anonymous):

can you please explain it to me how you got that

OpenStudy (anonymous):

i got that 8/20 is for black ball

OpenStudy (anonymous):

total = 20 black = 8 first event to choose any black from 20 balls = 8/20 second event the same , but now we have only 19 balls remaining so 7/19 third event , now only 18 balls remaining and also 6 black remaining so 6/18.

OpenStudy (anonymous):

but over here we choosing 3 black balls

OpenStudy (anonymous):

does it means 1 black and 2 random. actually that's the part that i didn;t get

OpenStudy (anonymous):

@BangkokGarrett will help ya... he's good in explaining probabilities.

OpenStudy (anonymous):

thanks and what would be the answer for e

OpenStudy (anonymous):

could you please show me the answer for that E as well please

OpenStudy (anonymous):

0.154 i guess

OpenStudy (anonymous):

8/20 * 12/19 * 11/18 12 = all balls except black ones

OpenStudy (anonymous):

maybe its right i'm not sure

OpenStudy (anonymous):

i 'll tell u tommorrow, ok?

OpenStudy (anonymous):

ok can you please check a answer me for a question?

OpenStudy (anonymous):

A process is thought to output an overall rate of 4% nonconforming parts. In a random sample of 50 such parts, what is the probability of 2 nonconforming parts? Solve the problem to four significant digits, in two ways-using the binomial distribution and the Poisson approximation to the binomial distribution. for this question i used binomial distribution ans i got the answer 0.2762

OpenStudy (anonymous):

and for this question do i have to solve it by both methods?

OpenStudy (anonymous):

that's not of my level.. sorry can't help with this one.

OpenStudy (anonymous):

ok no problem. could you help me with this question Determine the probability of the following: a) An odd number appears in a single toss of a fair die. Answer i got: 0.5 b) At least one tail appears in two tosses of a coin. Answer i got: 1/4 = 0.25 c) The sum 7 appears in a single toss of a pair of fair die. need help with this one

OpenStudy (anonymous):

total outcomes 6^2 = 36 Now we have to calculate the numbers that make 7. for example: 1+6 =7 2+5 = 7 3+4=7 and there opposite 6+!, 5+2, 4+3 total = 6 P = 6 / 36 = 1/6 yer ans??

OpenStudy (anonymous):

ohhh i got it thanks

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks In a country club, 80% of women play tennis, 20% play pool and 10% play both tennis and pool. If a woman is chosen at random, find the probability: a) that she plays tennis or pool. b) that she plays neither tennis or pool. c) if she plays tennis what is the probability she plays pool? d) if she plays pool what is the probability she play tennis?

OpenStudy (anonymous):

a) answer is 90% b) answer is 10% c) 10%/80% = 1/8 d) 10%/20% = 0.5

OpenStudy (anonymous):

these are the answers that i got

OpenStudy (anonymous):

c and d are correct... dunno bout a and b

OpenStudy (anonymous):

thank you so much for helping today, i really appreciate it. :)

OpenStudy (anonymous):

yw..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8 can u explain me part D?

ganeshie8 (ganeshie8):

In a country club, 80% of women play tennis, 20% play pool and 10% play both tennis and pool. If a woman is chosen at random, find the probability: a) that she plays tennis or pool. b) that she plays neither tennis or pool. c) if she plays tennis what is the probability she plays pool? d) if she plays pool what is the probability she play tennis?

ganeshie8 (ganeshie8):

is this the question ?

OpenStudy (anonymous):

no,, question related to balls.

OpenStudy (anonymous):

A ball is drawn from a box containing 5 red, 8 black and 7 blue balls. Determine the probability that: a) if a ball is drawn at random, it is black. answer i got: 0.40 b) if two balls are drawn, they are both black. answer i got: 0.1474 c) if three balls are drawn, exactly one is black, one is blue and one is red. d) if three balls are drawn, at least one is black. e) if four balls are drawn, at most one of them is black. Note: All selection is without replacement. can you please check the answers i got is right or wrong and can you help me solve c, d, and e too.

OpenStudy (anonymous):

this one

ganeshie8 (ganeshie8):

d) if three balls are drawn, at least one is black.

ganeshie8 (ganeshie8):

probability for having "at least one black" = 1 - probability for having NONE of them black

OpenStudy (anonymous):

P for NONE black = 12/20 * 11/19 ??? right??

ganeshie8 (ganeshie8):

yes, since it is without replacement, P for NONE of them black = 12/20*11/19*10/18

ganeshie8 (ganeshie8):

simply subtract it from 1, u wud get the probability for complement event : "atleast one of them black"

OpenStudy (anonymous):

11/57 ??

ganeshie8 (ganeshie8):

11/57 is for NONE of them black

OpenStudy (anonymous):

1 - 11/57 = 46/57

ganeshie8 (ganeshie8):

yess

OpenStudy (anonymous):

thanks... now part E. plz

OpenStudy (anonymous):

part E for exact one black ball

ganeshie8 (ganeshie8):

e) if four balls are drawn, at most one of them is black.

ganeshie8 (ganeshie8):

P for "atmost one of them black" = P for "NONE of them black" + P for "ONE of them black"

OpenStudy (anonymous):

we have calculated NONE, 11/57

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that was for three balls

ganeshie8 (ganeshie8):

yahh we need to calculate again

OpenStudy (anonymous):

12/20 * 11/19 * 10/18 * 9/17 = 11/ 57 * 9/17 = 33/323 right??

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

P for one black = 8/20 ???

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

8/20 * 7/19 * 6/18 * 5/17 ?????

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

remember you want "P for EXACTLY one of them black"

ganeshie8 (ganeshie8):

since you're drawing FOUR balls, first can be BLACK, and remaining needs to be other colors

ganeshie8 (ganeshie8):

second can be BLACK, and remaining needs to be other colors

ganeshie8 (ganeshie8):

third can be BLACK, and remaining needs to be other colors

ganeshie8 (ganeshie8):

fourth can be BLACK, and remaining needs to be other colors

OpenStudy (anonymous):

12/20 * 11/19 * 10/18 * 8/17???

ganeshie8 (ganeshie8):

nope, u need to consider FOUR cases and add all the probabilities

ganeshie8 (ganeshie8):

P(EXACTLY one ball black) = P(first ball black, all others not black) + P(second ball black, all others not black) + P(third ball black, all others not black) + P(fourth ball black, all others not black)

ganeshie8 (ganeshie8):

getting ?

ganeshie8 (ganeshie8):

four cases for getting exactly one black :- 1) \(BLACK, OTHER, OTHER, OTHER\) 2) \(OTHER, BLACK, OTHER, OTHER\) 3) \(OTHER, OTHER, BLACK, OTHER\) 4) \(OTHER, OTHER, OTHER, BLACK\)

ganeshie8 (ganeshie8):

when u work each case and add the probabilities u would get :- P(exactly one black ball) = \(\large \frac{8}{20}*\frac{12}{19} * \frac{11}{18} *\frac{10}{17} + \\ \frac{12}{20}*\frac{8}{19} * \frac{11}{18} *\frac{10}{17} + \\ \frac{12}{20}*\frac{11}{19} * \frac{8}{18} *\frac{10}{17} + \\ \frac{12}{20}*\frac{11}{19} * \frac{10}{18} *\frac{8}{17} \\ = \frac{352}{969} \)

OpenStudy (anonymous):

@ganeshie8 got it... thanks!!